OpenStudy (anonymous):

simple this question? I'm not good with fractions at all! 1 over 5(10a-4)+1 over 2(8 + 4a)

OpenStudy (anonymous):

\[\frac{ 1 }{ 5(10a-4)} + \frac{ 1 }{ 2(8+4a) }\]In order to add these, the denominators must be the same. So simply multiply the denominators here to get the common denominator you need to add them: \[\frac{ 1 }{ 50a-20 } + \frac{ 1 }{ 8a+16 } = \frac{ 1+1 }{ (50a-20)(8a+16) }\]Simplify: \[\frac{ 2 }{ (400a ^{2}+800a-160a-320) } = \frac{ 2 }{ 400a ^{2} + 640a -320 }\]Now just simplify further by using 2 as the GCF (greatest common factor): \[= \frac{ 1 }{ 200a ^{2}+320a-160 }\]

OpenStudy (anonymous):

thank you!:)